OpenStudy (muzzack):

The speed of a tsunami s (in meters/second) is approximated by s=sqrt of gd, where g is the acceleration due to gravity 9.8 m/s^2 and d is the depth of the water (in meters). If the speed of a tsunami is 200 m/s, about how deep is the water?

OpenStudy (muzzack):

jst need explanation not answer

OpenStudy (perl):

Basically this problem is what's known as 'plug and chug' . Plug in whats given into the equation given, and solve for the unknown.

OpenStudy (muzzack):

The speed of a tsunami s (in meters/second) is approximated by s=\(sqrt{gd}\) , where g is the acceleration due to gravity 9.8 \( m/s^2\) and d is the depth of the water (in meters). If the speed of a tsunami is 200 m/s, about how deep is the water?

OpenStudy (perl):

the equation is : s=sqrt(gd) the speed is given as 200 m/s we know from physics that g is about 9.81 m/s^2 200 = sqrt( 9.81* d ) solve for d the unknown

OpenStudy (muzzack):

\[\sqrt{gd}\]

OpenStudy (muzzack):

so \[200=(\sqrt{9.8}*d)\] = so you do the inverse of you graph the function to know what d is ? @perl

OpenStudy (muzzack):

or*

OpenStudy (muzzack):

@ganeshie8 ?